OpenStudy (anonymous):

A rectangular box is to be made to the following requirements: - The length must be one and a half times the width - The twelve edges must have a total length of 6m Find the dimensions of the box that meets these requirements and that maximizes the volume.

OpenStudy (anonymous):

Please help, using differentiation

OpenStudy (anonymous):

your width=w=0.4meters = 0.4m height h=0.5m length L=1.5w=1.5x.4=0.6 volume v=Lwh=0.6(4)(.5)=3/25= 0.12meter cube

OpenStudy (anonymous):

do you have the answer of choices there fetzer?

OpenStudy (anonymous):

Those are exactly correct, would you be able to explain as to how you obtained those answers :). Thank you very much

OpenStudy (anonymous):

ok bear with me fetzer..lol what ststes are you in? ill explain them very clearly...

OpenStudy (anonymous):

Ststes? do you mean States? I'm in Western Australia.

OpenStudy (anonymous):

ah ok,,,,

OpenStudy (anonymous):

draw there a rectangle in your scratch paper

OpenStudy (anonymous):

Yeah, don't wait for me just keep going :p

OpenStudy (anonymous):

ok sorry man..lol..

OpenStudy (anonymous):

you have width=w length = 1.5w perimeter P=4w+6w+4h P=10w +4h but it is given as 6meters P=10w +4h =6meter 4h =6 - 10w h=(3-5w)/2 now the volume V=Lwh V=1.5w(w)(h) V= 1.5w^2(h)

OpenStudy (anonymous):

now plug in your h=(3-5w)/2 to the volume V= 1.5w^2(h)

OpenStudy (anonymous):

Okay I am still with you :)

OpenStudy (anonymous):

V= 1.5w^2[(3-5w)/2] V=[(9/2)w^2 -(15/2w^3)]/2 V=[(9/2)w^2 -(15/2)w^3)]/2 V=[9w^2 -15w^3]/4

OpenStudy (anonymous):

factored out 3/4 there

OpenStudy (anonymous):

now you can do derivative of V with respect to w

OpenStudy (anonymous):

now solve for w

OpenStudy (anonymous):

V=(3/4)(3w^2-5w^3) dV/dw =(3/4)(6w-15w^2)=0 we zero them to get the maximum volume. 15w^2=6w 15w = 6 w=6/15=0.4

OpenStudy (anonymous):

Now L=1.5w=1.5(.4)=0.6m h=(3-5w)/2 =(3-5(.4))/2 =0.5 V= Lwh=.6(.5)(.4)=3/25=0.12m

OpenStudy (anonymous):

im sorry V=0.12meters cube

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

yw fetzer..good luck now....

OpenStudy (anonymous):

fetzerc, A solution using Mathematica is presented in the attachment. Unfortunately a solution as presented in the attachment cannot be produced in a couple of minutes.

OpenStudy (anonymous):

How do you use Mathematica

OpenStudy (anonymous):

Mathematica 8 is a very large (3Gb installation file) and comprehensive mathematics computer program able to generate symbolic as well as nummeric solutions. The program is not free, but the price for active students is very low. More about Mathematica can be found at: http://www.wolfram.com/solutions/education/students/ As a math hobbyist I own a Home edition of Mathematica 8 and have no financial interests in the Wolfram organization.

OpenStudy (anonymous):

A friend actually has a student copy and I'm wanting to useit, could you give a tutorial on how I would use it for the question I asked :P?

OpenStudy (anonymous):

Yes I guess I could help out. Mathematica is a very large system and the help file(s) is about 1Gb in size. The student copy should have a help system as part of it. I will have to admit that the thing is very intimidating at first.